Compute the resonant frequency of 31P, 1H, and 2H at 2.0 Tesla and at 3.0 Tesla?
Answer #1

Gyromagnetic ratio (γ) in MHz/T

γ_1H (Proton) = 42.58

γ_2H (Deutron) = 6.53

γ_31P (Phosphorus) = 17.24

Resonant Frequency, f_o = γB MHz

B1 = 2T

f_o_1H = 42.58*2 = 85.16MHz

f_o_2H = 6.53*2 = 13.06MHz

f_o_31P = 17.24*2 = 34.48MHz

B2 = 3T

f_o_1H = 42.58*3 = 127.74MHz

f_o_2H = 6.53*3 = 19.59MHz

f_o_31P = 17.24*3 = 51.72MHz